OpenStudy (cmkc109):

Confusion about transformation: Why do y=3^-x and y= (1/3 )^x have the same curve?

OpenStudy (tkhunny):

Exponent rules. \(3^{-x} = \dfrac{1}{3^{x}} = \left(\dfrac{1}{3}\right)^{x}\) Many times, the answer is not obvious. A little exploration is required.

OpenStudy (irishboy123):

\(y=3^{-x}\) \(\ln y=-x \ln 3\) different system \( y'= (1/3 )^{x'} \) \(\ln y'= x' ln (1/3 ) = - x' \ln (3 )\) wot he said :)

OpenStudy (cmkc109):

Thank you both of you !!!